EFFECTIVE RATE OF INTEREST

Find the interest rates earned on each of the following. Round each answer to two decimal places.

  1. You borrow $700 and promise to pay back $770 at the end of 1 year.

    %

  2. You lend $700 and the borrower promises to pay you $770 at the end of 1 year.

    %

  3. You borrow $93,000 and promise to pay back $279,116 at the end of 7 years.

    %

  4. You borrow $11,000 and promise to make payments of $3,359.50 at the end of each year for 5 years.

    %
Answer #1

a. r = 770/700 - 1

r = 0.1

r = 10%

You pay interest at the rate of 10% per year.

b. r = 770/700 - 1

r = 0.1

r = 10%

You earn interest at the rate of 10% per year.

c.  

d) N = 5

PMT = -3,359.50

PV = 11,000

FV = 0

Compute I/Y

I/Y = 15.99%

You can also compute r by solving the annuity equation.

Can you please upvote? Thank You :-)